REU: SCHEDULE FOR WEEKS FIVE AND SIX

Date  YSP  Kottwitz Constantin Rothenberg
Monday, July 15  9:30 – 2:30 4:30 – 6:00 2:45 – 4:15
Tuesday, July 16 9:30 – 2:30 2:45 – 3:45
Wednesday, July 17 9:30 – 2:30 4:15 – 5:45  2:45 – 3:45
Thursday, July 18 9:30 – 2:30  2:45 – 4:15
Friday, July 19  9:30 – 2:30 4:15 – 5:45  2:45 – 3:45
Monday, July 22  9:30 – 2:30 4:15 – 5:45  2:45 – 3:45
Tuesday, July 23  9:30 – 2:30  2:45 – 4:15
Wednesday, July 24 9:30 – 2:30 4:15 – 5:45  2:45 – 3:45
Thursday, July 25 9:30 – 2:30  2:45 – 4:15
Friday, July 26 9:30 – 2:00* 4:15 – 5:45 2:45 – 3:45

*Staff meeting 12:00 – 2:00

Graduate students will work with undergraduates during part of Kottwitz’s time slots, but extra problem sessions are not planned.

Graduate students in geometry plan to hold problem and discussion sessions; Tuesdays and Thursdays from 4:15 to 5:45 will probably be the regular times.

Quantum mechanics problem sessions will be organized on Monday.

Here is the schedule in summary form:

Fifth week:

YSP: 9:30 – 2:30 M – F Kottwitz 4:30 – 6:00 Monday; 4:15 – 5:45 Wednesday and Friday Constantin 2:45 – 4:15 Monday and Thursday Rothenberg 2:45 – 3:45 Tuesday, Wednesday, and Friday

Sixth week:

YSP 9:30 – 2:30 M – F (to 2:00 on Friday) Kottwitz 4:15 – 5:45 MWF Constantin 2:45 – 4:15 TTh Rothenberg 2:45 – 3:45 MWF